# Self-Assembled Chains and Solids of Dipolar Atoms in a Multilayer

### v1

@@ Abstract
We predict that ultracold bosonic dipolar gases, confined within a multilayer geometry, may undergo self-assembling processes, leading to the formation of chain gases and solids. These dipolar chains, with dipoles aligned across different layers, emerge at low densities and resemble phases observed in liquid crystals, such as nematic and smectic phases. We calculate the phase diagram using quantum Monte Carlo methods, introducing a newly devised trial wave function designed for describing the chain gas, where dipoles from different layers form chains without in-plane long-range order. We find gas, solid, and chain phases, along with quantum phase transitions between these states. Specifically, we predict a quantum phase transition from a gaseous to a self-ordered phase, which occurs at a critical interlayer distance. Remarkably, in the self-organized phases, the mean interparticle distance can significantly exceed the characteristic length of the interaction potential, yielding solids and chain gases with densities several orders of magnitude lower than those of conventional quantum solids.
@@
Recent experiments with ultracold atomic gases have achieved significant advances in the observation of self-organized systems, such as solids with superfluid properties (supersolids) observed in platforms with spin-orbit coupling Li et al. (2017) and long-range cavity-mediated interactions Léonard et al. (2017). Although in those setups spatial translational symmetry is broken, actually the crystal periodicity is externally imposed either by optical lattices or cavities. This limitation was recently surpassed in experiments with dipolar quantum gases, which reported a spontaneous formation of crystals of droplets Chomaz et al. (2019); Böttcher et al. (2019); Tanzi et al. (2019). The realization of a quantum self-assembled system, in which continuous translational symmetry is naturally broken, remains an open challenge within the community of ultradilute gases.
@@
A relevant question that arises is if stable quantum solids and chains can be spontaneously formed at ultradilute densities, where the mean interparticle distance is much larger than the range of the potential? This scenario is unusual in condensed matter, where solid phases typically occur at high densities (a notable exception is that of Wigner crystals Wigner (1934) formed at low density, however, quantum effects are typically small). While there have been studies on low-density dipolar chains, previous research either neglected interactions between chains Wang et al. (2006) or restricted themselves to the regime of small spatial fluctuations, about the crystal phase You and Wang (2014). As a result, complete description of self-assembled systems at ultradilute densities is still missing.
In the present work, we propose a possible physical realization of quantum self-assembled states. The platform consists of a system of dipolar bosons confined to a multilayer geometry, as schematically shown in Fig. 1. Using exact many-body quantum Monte Carlo techniques Hammond et al. (1994); Toulouse et al. (2015); Boronat and Casulleras (1994) and proposing a novel trial wave function tailored specifically to model chain gases, we determine the ground-state phase diagram when the dipoles are all aligned perpendicularly to the parallel layers by an external field. We anticipate the emergence of a stable phase characterized by self-organized chains across layers when the density is low. We find a quantum phase transition from a gas to an ordered phase, solid or chains, as the interlayer distance decreases. Remarkably, the density of the self-assembled phases is several orders of magnitude lower than the one of conventional quantum solids.
Figure 2: Ground-state phase diagram of dipolar bosons confined to a four-layer system at zero temperature, as a function of the total density $nr_{0}^{2}$ and the separation between layers $h/r_{0}$. We have identified three phases: gas, solid, and chains. In the limit of $h/r_{0}\to 0$, the low-density chain phase (indicated with a red bar) experiences a phase transition to a solid state (indicated with a blue bar) at the critical density $n_{\rm c}r_{0}^{2}\approx 0.283$.
Hamiltonian. We consider $N$ bosons of mass $m$ and dipole moment $d$ constrained by a multilayer geometry composed of $M$ two-dimensional (2D) parallel layers, each separated by a distance $h$. We assume that each layer contains the same number ($N/M$) of dipoles and that the interlayer tunneling is suppressed so that the transverse degrees of freedom are frozen. The dipolar moments are considered to be oriented perpendicularly to the 2D planes, which can be achieved by applying a polarizing field. The Hamiltonian of this system is given by the sum of the kinetic energy operator and the dipolar interaction potential,
@@
Quantum Monte Carlo Simulations. The ground-state properties of the system, described by the Hamiltonian (1), are calculated using the diffusion Monte Carlo (DMC) method Hammond et al. (1994); Toulouse et al. (2015); Boronat and Casulleras (1994). The DMC method has been shown to give an accurate description of correlated quantum systems Ceperley (2010). The DMC method solves stochastically the many-body Schrödinger equation in imaginary time resulting in an exact value (within controllable statistical error) of the ground-state energy, alongside other properties. We use guiding wave functions to impose a specific symmetry to the state and to improve the convergence. To characterize the gas phase, we use a trial wave function in Bijl-Jastrow (BJ) form, which maintains translational symmetry and consists of two-body correlation terms. Conversely, for the solid phase, we adopt a Nosanow-Jastrow (NJ) wave function. This NJ form incorporates one- and two-body correlation factors, thus breaking translational symmetry (see Supplemental Material (SM)). The positions of the dipoles confined to the multilayer geometry are projected into a box of size $L_{x}\times L_{y}$ and periodic boundary conditions along $(x,y)$ plane are used to simulate the bulk system. The total density is defined as $n=N/(L_{x}L_{y})$.
In addition to the gas and solid phases, the presence of anisotropic interactions allows for the emergence of a dipolar chain phase. This phase is similar to the molecular phase observed in bilayer systems Macia et al. (2014); Astrakharchik et al. (2016); Guijarro et al. (2022) but extends beyond two layers. While the use of BJ (for gas) and NJ (for crystal) trial functions is well-established in the literature, to the best of our knowledge, a Monte Carlo description for dipolar chain systems is still lacking. Here, we propose a new type of trial wave function specifically designed to describe chain phases. Its primary characteristic lies in not only capturing interlayer correlations but also enabling the free movement of chains within planes. We found that the chain trial wave function results in lower energy in certain regions of the phase diagram compared to the trial wave functions for solids and gases. Therefore, the novel introduction of the chain trial wave function is essential for fully describing the phase diagram of the dipolar multilayer system (see SM).
@@
Ground state of dipolar bosons within a four-layer system. To construct the phase diagram, we explored the parameter space and calculated the ground-state energy using three different trial wave functions. To verify which phase has the lowest energy, we calculate the DMC energy $E_{\rm DMC}$, add the tail energy correction $E_{\rm tail}$, and extrapolate the corrected value to the thermodynamic limit, $E_{\rm th}$, as discussed in the SM. The resulting phase diagram is reported in Fig. 2 as a function of two key parameters: the dimensionless density $nr_{0}^{2}$ and the separation between layers $h/r_{0}$. An important observation is the existence of the chain phase which spans across a wide range of interlayer distances. Its large extension enhances the feasibility of its experimental observation. Furthermore, this phase extends to extremely low densities, at which the mean interparticle distance within one layer can range from 1 to 20 times larger than the dipolar length $r_{0}$, resulting in the formation of atomic dipolar chains at ultradilute density.
At small interlayer distances $h/r_{0}<0.2$, we observe the appearance of the solid phase as the density increases, above the crystallization density of approximately $n_{\rm c}r_{0}^{2}\approx 0.283$. Additionally, for interlayer distances larger than $h/r_{0}>0.4$, we observe a quantum phase transition from chains to gas as the density increases. Remarkably, for certain densities ($nr_{0}^{2}\approx 1$), the change in $h/r_{0}$ induces a series of phase transitions crossing gas, chain, and solid phases as the distance between the layers is decreased. We provide estimates for the boundaries of the gas phase, along with the transition line from chains to gas. However, determining the precise location of the transition from chains to solid is computationally challenging due to the similar energies of these phases across a broad parameter range.
@@
Various phases exhibit significantly different spatial distribution of dipoles, as quantified by the pair distribution function $g(r,l)$, which indicates the probability of finding two atoms at a relative distance $r$ and separated by $l$ layers. In Fig. 3, we show characteristic examples of the pair distribution functions $g(r,l)$ in the gas (a), solid (b), and chain (c) phases.
@@
Figure 3: Top panel: Typical in-plane $g(r,l=0)$ and off-plane $g(r,l>0)$ pair distribution functions in the gas (a), solid (b), and chain (c) phases within a four-layer geometry. Bottom panel: Characteristic snapshots of the atom coordinates from DMC simulations in the gas (d), solid (e), and chain (f) phases. The parameters are: gas $nr_{0}^{2}=2.0$, $h/r_{0}=0.5$; solid $nr_{0}^{2}=4.0$, $h/r_{0}=0.3$; chain $nr_{0}^{2}=0.01$, $h/r_{0}=0.9$.
@@
The same-layer distribution $g(r,l=0)$ vanishes at short distances due to the strongly repulsive hard core of the dipolar potential at $r/r_{0}\to 0$. In the gas phase, the in-layer correlation function exhibits a shallow maximum and then approaches the unit value, i.e., the asymptotic value of uncorrelated atoms, see Fig. 3a. However, correlation functions between different layers ($l>0$) show a notably distinct short-range behavior. In particular, there is a high peak in $g(r,l>0)$ at $r/r_{0}=0$ as a result of the maximum interlayer attraction between the dipoles in adjacent layers. As the distance $r$ increases, $g(r,l>0)$ exhibits a minimum, and eventually approaches its uncorrelated value of $1$. A representative snapshot of the atom positions in the gas phase in the DMC simulations is presented in Fig. 3d and it shows that the dipoles are uniformly distributed without any specific pattern.
@@
Characteristic examples of the pair distribution functions in the solid phase are also shown in Fig. 3b. At short distances, $r\to 0$, the behavior is similar to that observed in the gas phase, where the same-layer distribution vanishes as dipoles approach each other, while different-layer distributions exhibit a pronounced correlation peak. As distance $r$ increases, the solid pair distributions show stronger correlations compared to the gas phase, as a result of the spatial ordering in the solid. Consequently, $g(r,l)$ displays several pronounced oscillations in a solid phase, whereas in the gas phase, it rapidly converges to the uncorrelated values of $g(r,l)=1$. The amplitude of these oscillations depends on the system parameters. In a triangular lattice, the first peak is particularly prominent, corresponding to six neighbors located at the lattice spacing $r\approx a_{latt}$, which is proportional to the mean interparticle distance $a_{latt}=(2/\sqrt{3})^{1/2}r_{m}\approx 1.075r_{m}$ and is determined by the in-layer density, $r_{m}=\sqrt{M/n}$. In Fig. 3e, a snapshot of the atom coordinates during the DMC simulation of the solid phase is presented, revealing the periodic structure of a triangular lattice.
@@
In Fig. 3c, we report the pair distribution functions $g(r,l)$ for the chain phase. We observe intermediate features, where some aspects resemble those of a gas phase while others are closer to those of a crystal phase. Specifically, the in-plane pair distribution $g(r,l=0)$ shows correlations typical of the gas phase, indicating no long-range order within the layer. However, the off-plane pair distributions $g(r,l>0)$ exhibit high peaks at short distances, similar to the strong inter-layer correlations observed in the solid phase. The snapshots of the atom coordinates during the DMC simulation, shown in Fig. 3f, provide insight into the structure of this unconventional phase. Here, the dipolar chains are uniformly distributed without forming a regular structure, indicating the absence of positional order. However, the dipoles forming the chains exhibit strong correlations in the direction perpendicular to the planes. This behavior resembles that of a nematic phase in liquid crystals.
@@
Quantum phase transition: from gas to a self-ordered phase. The effect of different number of layers (from one to ten) on the phase diagram is summarized in Fig. 4, where the schematic phase diagram of the multilayer system is shown as a function of the dimensionless density $nr_{0}^{2}$ and the separation between layers $h/r_{0}$. The thick horizontal lines for small $h/r_{0}$ correspond to the critical freezing density $n_{\rm c}r_{0}^{2}=290/M^{5}$, indicating the density at which an $M$-layer system of dipolar bosons will form a crystal of super-dipoles (where all dipoles align perfectly at each point) when the interlayer separations are extremely small $h\to 0$ (see SM). Notably, this estimation of the freezing point shows that the freezing density decreases as $1/M^{5}$ with the number of planes $M$. This strong dependence on $M$ suggests that the multilayer geometry offers a very promising setup for observing quantum solids with low surface densities in future experiments. This finding is significant because increasing the number of planes in the system can yield a very dilute solid, especially in situations characterized by small interlayer distances.
The phase diagram shown in Fig. 4 also reports an approximate threshold density $n_{\rm t}r_{0}^{2}={M}/[{2(M-1)^{2}h^{2}}]$ (shown with dashed lines) as a function of the interlayer separation for different number of planes (as discussed in the SM). The position of the threshold, $n_{\rm t}r_{0}^{2}$, is defined as the density at which the classical interaction between particles in the top and bottom layers changes its sign at the mean interparticle distance. Densities larger (smaller) than this threshold favor the formation of a gas (self-ordered) phase. According to this estimate, a self-ordered phase can occur at very low densities, not only for small interlayer distances but also across a wide range of separations. To support these crude estimates, we performed DMC simulations.
@@
Using the DMC method, we observed a quantum phase transition from a gas to an ordered phase (solid or chains) as the interlayer distance decreased. To determine the critical interlayer values for this transition, we fixed the number of planes $M$ and the total density $nr_{0}^{2}$, and then systematically reduced the interlayer distance while keeping the density constant. We calculated the ground-state energy using both gas and ordered (solid or chains) trial wave functions. The phase at each point corresponds to the phase with the lowest energy. The critical interlayer distance corresponds to the value at which the phase changes from gas to an ordered phase. The estimated critical interlayer distances for the gas-ordered transition are reported in the phase diagram of Fig. 4, covering the range of layers from $M=2$ up to 10. For each $M$ value, the critical interlayer distance is reported for one density. One might notice that the simple approximation of the threshold density $n_{\rm t}r_{0}^{2}$ becomes less accurate as $M$ increases. Remarkably, the self-ordered phase exists across a broad range of densities and interlayer separations. Adjusting the number of planes can alter the density of the phase significantly, decreasing from $nr_{0}^{2}=10$ to $0.001$ as the number of layers increases from $M=2$ to 10. These extremely low densities correspond to mean interparticle distances ranging from $0.4$ to 100 times the dipolar length $r_{0}$. Moreover, the range of interlayer distances where the self-assembled phase is stable expands with the number of layers, which suggests the possibility of its experimental realization.
@@
Figure 4: Phase diagram for various number of layers ($M=1,\cdots,M=10$) as a function of the total density $nr_{0}^{2}$ and the separation between layers $h/r_{0}$ Quantum phase transition from a gas to a self-ordered phase occurs as the interlayer distance decreases. The squares indicate DMC predictions for the gas-ordered phase transitions, considering a single value of density $nr_{0}^{2}$ for each number of layers $M$ and changing interlayer separation $h$. The crystallization $n_{\rm c}r_{0}^{2}$ and threshold $n_{\rm t}r_{0}^{2}$ densities are shown by the thick and dashed lines, respectively.
@@
Conclusions. We used the diffusion Monte Carlo method to obtain the phase diagram of two-dimensional dipolar bosons confined within a multilayer geometry at zero temperature. Our findings indicate the possibility of generating self-assembled phases, such as solid and chains, at ultra-low densities. Remarkably, the density of these systems can be several orders of magnitude lower than that of conventional solids. Moreover, the freezing density can be further decreased by increasing the number of layers. Our results present a novel example of quantum dipolar self-ordered phases characterized by ultradilute density. In the solid phase, unlike previous realizations, the dipolar interactions by themselves are sufficient to create the crystalline structure, which consists of individual atoms rather than clusters. Whether this crystal exhibits a finite superfluid signal and qualifies as a new supersolid remains a topic for future research.

### v2

@@ Abstract
We predict that ultracold bosonic dipolar gases, confined within a multilayer geometry, may undergo self-assembling processes, leading to the formation of chain gases and solids. These dipolar chains, with dipoles aligned across different layers, emerge at low densities and resemble phases observed in liquid crystals, such as nematic and smectic phases. We calculate the phase diagram using quantum Monte Carlo methods, introducing a newly devised trial wave function designed for describing the chain gas, where dipoles from different layers form chains without in-plane long-range order. We find gas, solid, and chain phases, along with quantum phase transitions between these states. Specifically, we predict the existence of quantum phase transitions from gaseous to self-ordered phases, as the interlayer distance is decreased. Remarkably, in the self-organized phases, the mean interparticle distance can significantly exceed the characteristic length of the interaction potential, yielding solids and chain gases with densities several orders of magnitude lower than those of conventional quantum solids.
@@
Recent experiments with ultracold atomic gases have achieved significant advances in the observation of self-organized systems, such as superfluid stripes reported in platforms with spin-orbit interactions Li et al. (2017); Geier et al. (2023). Additionally, solids with superfluid properties (supersolids) have been observed in setups with long-range cavity-mediated interactions Léonard et al. (2017). In these latter setups, although spatial translational symmetry is broken, the crystal periodicity is externally imposed by cavities. This limitation was recently surpassed in experiments with dipolar quantum gases, which reported a spontaneous formation of crystals of droplets Chomaz et al. (2019); Böttcher et al. (2019); Tanzi et al. (2019). The realization of an atomic quantum self-assembled system, in which continuous translational symmetry is naturally broken, remains an open challenge within the community of ultradilute gases.
@@
A relevant question that arises is if stable quantum solids and chains can be spontaneously formed at ultradilute densities, where the mean interparticle distance is much larger than the range of the potential. This scenario is unusual in condensed matter, where solid phases typically occur at high densities (a notable exception is that of Wigner crystals Wigner (1934) formed at low density; however, quantum effects are typically small). While there have been studies on low-density dipolar chains, previous research either neglected interactions between chains Wang et al. (2006) or restricted themselves to the regime of small spatial fluctuations, about the crystal phase You and Wang (2014). As a result, a complete description of self-assembled systems at ultradilute densities is still missing.
In the present article, we propose a possible physical realization of quantum self-assembled states. The platform consists of a system of dipolar bosons confined to a multilayer geometry, as schematically shown in Fig. 1. Using exact many-body quantum Monte Carlo techniques Hammond et al. (1994); Toulouse et al. (2015); Boronat and Casulleras (1994) and proposing a novel trial wave function tailored specifically to model chain gases, we determine the ground-state phase diagram when the dipoles are all aligned perpendicularly to the parallel layers by an external field. We anticipate the emergence of a stable phase characterized by self-organized chains across layers when the density is low. We find a quantum phase transition from a gas to an ordered phase, solid or chains, as the interlayer distance decreases. Remarkably, the density of the self-assembled phases is several orders of magnitude lower than the one of conventional quantum solids.
Figure 2: Ground-state phase diagram of dipolar bosons in a four-layer system at zero temperature as a function of the total density $nr_{0}^{2}$ and the interlayer distance $h/r_{0}$. Symbols indicate the phases according to the DMC calculations: triangles for chains, squares for solid, and dots for gas. Where solid and chain phases overlap, two-colored squares are used. The coloring qualitatively shows the extension of each phase. As $h/r_{0}\to 0$, the low-density chain phase (red bar) transitions to a solid state (blue bar) at the critical density $n_{\rm c}r_{0}^{2}\approx 0.283$.
Hamiltonian. We consider $N$ bosons of mass $m$ and dipole moment $d$ constrained by a multilayer geometry composed of $M$ two-dimensional (2D) parallel layers, each separated by a distance $h$. We assume that each layer contains the same number ($N/M$) of dipoles and that the interlayer tunneling is suppressed so that the transverse degrees of freedom are frozen. The dipolar moments are considered to be oriented perpendicularly to the 2D planes, which can be achieved by applying a polarizing field. The Hamiltonian of this system is given by the sum of the kinetic energy operator and the dipolar interaction potential:
@@
Quantum Monte Carlo simulations. The ground-state properties of the system, described by the Hamiltonian (1), are calculated using the diffusion Monte Carlo (DMC) method Hammond et al. (1994); Toulouse et al. (2015); Boronat and Casulleras (1994). This method has been shown to give an accurate description of correlated quantum systems Ceperley (2010). The DMC algorithm solves stochastically the many-body Schrödinger equation in imaginary time resulting in an exact value (within controllable statistical error) of the ground-state energy, alongside other properties. As usual in DMC, we use guiding wave functions for importance sampling to reduce the variance. We use three trial wave functions: for a gas, solid, and chains. The positions of the dipoles are projected into a box of size $L_{x}\times L_{y}$ with periodic boundary conditions along $(x,y)$. The box size is related to the total density according to $n=N/(L_{x}L_{y})$ (further details in the end matter section).
@@
Ground state of dipolar bosons within a four-layer system. To construct the phase diagram, we explored the parameter space (dimensionless density $nr_{0}^{2}$ and interlayer distance $h/r_{0}$) and calculated the ground-state energy using the three different trial wave functions. The phase at each point corresponds to the phase that yields the lowest energy. We calculate the DMC energy $E_{\rm DMC}$, add the tail energy correction $E_{\rm tail}$, and extrapolate the corrected value to the thermodynamic limit, $E_{\rm th}$, as discussed in Supplemental Material  [^1] See Supplemental Material, which includes Refs. Boronat and Casulleras (1994); Hammond et al. (1994); Toulouse et al. (2015); Ceperley (2010); Cikojević et al. (2019); Parisi et al. (2019); Comparin et al. (2019); Guijarro et al. (2020, 2022); Ardila et al. (2020); Bombín et al. (2019, 2017); Rodríguez-López et al. (2021); Sánchez-Baena et al. (2018); Reatto and Chester (1967); Cazorla and Boronat (2017); Guijarro (2020); Astrakharchik et al. (2007), for additional information about the trial wave functions, freezing density, and finite size effects.. The resulting phase diagram is reported in Fig. 2 as a function of two key parameters: $nr_{0}^{2}$ and $h/r_{0}$. An important observation is the existence of the chain phase which spans across a wide range of interlayer distances. Its large extension enhances the feasibility of its experimental observation. Furthermore, this phase extends to extremely low densities, at which the mean interparticle distance within one layer can range from 1 to 20 times larger than the dipolar length $r_{0}$, resulting in the formation of atomic dipolar chains at ultradilute density. For $h/r_{0}\lesssim 0.2$, the strong interlayer attraction between dipoles results in the formation of tightly bound chains of a small $(x-y)$ spatial size. As the density is increased, the strong in-plane dipolar repulsion pushes the system toward crystallization, which eventually occurs at the critical density $n_{\rm c}r_{0}^{2}\approx 0.283$. While for small $h$, the radius of the attractive cone is small, resulting in rigid chains, for larger $h$ the chains become flexible and loosely bound. Consequently, an increase in the density for large values of $h$ leads to a transition from chains to gas rather than to a solid. Remarkably, for certain densities ($nr_{0}^{2}\approx 1$), the decrease in $h/r_{0}$ induces a series of phase transitions crossing gas, chain, and solid phases. We provide estimates for the boundaries of the gas phase, along with the transition line from chains to gas. However, determining the precise transition from chains to solid is computationally challenging due to the similar energies of these phases across a broad parameter range.
@@
Various phases exhibit significantly different spatial distribution of dipoles, as quantified by the pair distribution function $g(r,l)$, which indicates the probability of finding two atoms at a relative in-plane distance $r$ and separated by $l$ layers. In Fig. 3, we show characteristic examples of the pair distribution functions $g(r,l)$ in the gas (a), solid (b), and chain (c) phases.
@@
Figure 3: Typical in-plane $g(r,l=0)$ and off-plane $g(r,l>0)$ pair distribution functions in the gas (a), solid (b), and chain (c) phases within a four-layer geometry. Inset: characteristic snapshots of the atom coordinates in the corresponding phase. The parameters are: gas, $nr_{0}^{2}=2.0$, $h/r_{0}=0.5$; solid, $nr_{0}^{2}=4.0$, $h/r_{0}=0.3$; and chain, $nr_{0}^{2}=0.01$, $h/r_{0}=0.9$.
@@
In the gas phase, the in-layer distribution ($l=0$) vanishes at short distances due to a strong dipolar repulsion for $r\to 0$. For larger distances, $g(r,l=0)$ exhibits a shallow maximum and then approaches the unit value, i.e., the asymptotic value of uncorrelated atoms; see Fig. 3a. However, correlation functions between different layers ($l>0$) show a notably distinct short-range behavior. In particular, there is a high peak in $g(r,l>0)$ due to strong dipolar attraction for $r\to 0$. For larger $r$, $g(r,l>0)$ exhibits a minimum, and eventually approaches its uncorrelated value of $1$. A representative snapshot of the atom positions in the gas phase in the DMC simulations is presented in the inset of Fig. 3a, showing that the dipoles are uniformly distributed without any specific pattern.
@@
Characteristic examples of the pair distribution functions in the solid phase are shown in Fig. 3b. At short distances, $r\to 0$, the behavior is similar to that observed in the gas phase, where the same-layer distribution vanishes as dipoles approach each other, while different-layer distributions exhibit a pronounced correlation peak. For larger distances, the pair distributions show stronger correlations in the solid phase as compared to the gas one, as a result of the spatial ordering in the solid. Consequently, $g(r,l)$ displays several pronounced oscillations in a solid phase, whereas in the gas phase, it rapidly converges to the uncorrelated values of $g(r,l)=1$. The amplitude of these oscillations depends on the system parameters. In a triangular lattice, the first peak is particularly prominent, corresponding to six neighbors located at the lattice spacing $r\approx a_{\text{latt}}$, which is proportional to the mean interparticle distance $a_{\text{latt}}=(2/\sqrt{3})^{1/2}r_{m}\approx 1.075r_{m}$ and is determined by the in-layer density, $r_{m}=\sqrt{M/n}$. In the inset of Fig. 3b, a snapshot of the atom coordinates during the DMC simulation of the solid phase is presented, revealing the periodic structure of a triangular lattice.
@@
In Fig. 3c, we report the pair distribution functions $g(r,l)$ for the chain phase. We observe intermediate features, where some aspects resemble those of a gas phase while others are closer to those of a crystal phase. Specifically, the in-plane pair distribution $g(r,l=0)$ shows correlations typical of the gas phase, indicating no long-range order within the layer. However, the off-plane pair distributions $g(r,l>0)$ exhibit high peaks at short distances, similar to the strong interlayer correlations observed in the solid phase. The snapshots of the atom coordinates during the DMC simulation, shown in the inset in Fig. 3c, provide insight into the structure of this unconventional phase. Here, the dipolar chains are uniformly distributed without forming a regular structure, indicating the absence of positional order. However, the dipoles forming the chains exhibit strong correlations in the direction perpendicular to the planes. This behavior resembles that of a nematic phase in liquid crystals.
@@
Quantum phase transition: From gas to a self-ordered phase. The effect of different numbers of layers (from one to ten) on the phase diagram is summarized in Fig. 4, where the schematic phase diagram of the multilayer system is shown as a function of the dimensionless density $nr_{0}^{2}$ and the separation between layers $h/r_{0}$. The thick horizontal lines for small $h/r_{0}$ correspond to the critical freezing density $n_{\rm c}r_{0}^{2}=290/M^{5}$, indicating the density at which an $M$-layer system of dipolar bosons will form a crystal of superdipoles (where all dipoles align perfectly at each point) when the interlayer separations are extremely small $h\to 0$ (see Supplemental Material Note (1)). Notably, this estimation of the freezing point shows that the freezing density decreases as $1/M^{5}$ with the number of planes $M$. This strong dependence on $M$ suggests that the multilayer geometry is a promising setup for observing quantum solids with low surface densities in future experiments. Increasing the number of planes can yield a very dilute solid, especially for small interlayer distances.
@@
Using the DMC method, we observed a quantum phase transition from a gas to an ordered phase (solid or chains) as the interlayer distance $h/r_{0}$ decreased. To determine the critical interlayer values for this transition, we fixed the number of planes $M$ and the total density $nr_{0}^{2}$, and then systematically reduced the interlayer distance $h/r_{0}$ while keeping the density constant. We calculated the ground-state energy using both gas and ordered (solid or chains) trial wave functions. The phase at each point corresponds to the phase with the lowest energy. The critical interlayer distance corresponds to the value at which the phase changes from gas to an ordered phase. The estimated critical interlayer distances for the gas-ordered transition are reported in the phase diagram in Fig. 4, covering the range from $M=2$ up to 10. For each $M$ value, the critical interlayer distance is reported for a single density, except in the case of $M=4$, where it is reported for multiple density values. Remarkably, the self-ordered phase exists across a broad range of densities and interlayer separations. Triangles indicate the gas-ordered phase transition for $M=4$ layers. At a fixed density, decreasing $h/r_{0}$ leads to a transition from gas to a self-ordered phase. This reduction in $h/r_{0}$ also diminishes the radius of the attractive cone of the dipolar interaction, aligning dipoles along the layers. Conversely, at a fixed $h/r_{0}$, decreasing the density induces to the transition. This is because lower density reduces in-plane repulsion, while interlayer attraction depends primarily on $h/r_{0}$. Overall, the transition density decreases as $h/r_{0}$ increases, and this behavior in the four-layer system is likely scalable to systems with more layers. Adjusting $M$ can significantly alter the density at which an ordered phase appears, decreasing from $nr_{0}^{2}=10$ to $0.001$ as $M$ increases from 2 to 10. Such extremely low densities correspond to mean interparticle distances ranging from $0.4$ to 100 times the dipolar length $r_{0}$. Moreover, the range of $h$ for stable self-assembled phases expands with $M$, which suggests the possibility of its experimental realization.
@@
Figure 4: Phase diagram for layers $M=1,\cdots,10$ as a function of the total density $nr_{0}^{2}$ and the interlayer distance $h/r_{0}$. A quantum phase transition from a gas to a self-ordered phase occurs as $h/r_{0}$ decreases. Squares indicate DMC predictions for the gas-ordered transitions at a single value of $nr_{0}^{2}$ while varying $h/r_{0}$ for each $M$. Triangles denote DMC results for $M=4$, varying both $nr_{0}^{2}$ and $h/r_{0}$. Thick lines show the crystallization densities $n_{\rm c}r_{0}^{2}$ for various $M$.
We propose that the $M$-dipole condensate, defined as the macroscopic eigenvalue of the $M$-body density matrix, could be serve as an order parameter to distinguish between different phases. The atomic condensate ($M=1$) is finite in the gas phase but vanishes in the chain and solid phases. In the bilayer case, it has been shown that the molecular condensate ($M=2$) becomes finite when chains of length $M=2$ are formed Macia et al. (2014). Similarly, we expect that when a chain of length $M$ is formed, the $M$-dipole condensate is present while all other condensates formed from $M-1,M-2,\cdots,1$ dipoles vanish. In contrast, condensation would be entirely absent in the solid phase.
@@
Conclusions. We used the diffusion Monte Carlo method to obtain the phase diagram of two-dimensional dipolar bosons confined within a multilayer geometry at zero temperature. Our findings indicate the possibility of generating self-assembled phases, such as solid and chains, at ultralow densities. Remarkably, the density of these systems can be several orders of magnitude lower than that of conventional solids. Moreover, the freezing density can be further decreased by increasing the number of layers. Our results present a novel example of quantum dipolar self-ordered phases characterized by ultradilute density. In the solid phase, unlike previous realizations, the dipolar interactions by themselves are sufficient to create the crystalline structure, which consists of individual atoms rather than clusters. Whether this crystal exhibits a finite superfluid signal and qualifies as a new supersolid remains a topic for future research.